# Non-stationary Gaussian random fields on hypersurfaces: Sampling and strong error analysis

### v1

@@ 3.3. Error analysis of the SFEM discretization
$\|\mathcal{Z}-\mathsf{Z}_{h}^{\ell}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}\lesssim C_{\alpha}(h)h^{\min\{\alpha-d/4;2\}}+|\log(h)|^{(d-1)/2}\leavevmode\nobreak\ h^{2}.$
Let then $\epsilon_{h}$ be the quantity defined by $\epsilon_{h}=\big{(}K_{\alpha}(h)h^{\min\{\alpha-d/4;2\}}\big{)}^{-1}\|\mathcal{Z}-\mathsf{Z}_{h}^{\ell}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}$, which satisfies
$\epsilon_{h}$ $\lesssim C_{\alpha}(h)K_{\alpha}(h)^{-1}+|\log(h)|^{(d-1)/2}K_{\alpha}(h)^{-1}h^{2-\min\{\alpha-d/4;2\}}.$
If $\alpha-d/4\in(0,2)$, we have $C_{\alpha}(h)=K_{\alpha}(h)=1$ and $02-\min\{\alpha-d/4;2\}>$ 0, and therefore $\epsilon_{h}\lesssim 1+$. Similarly, if $\alpha-d/4=2$, it holds $C_{\alpha}(h)=K_{\alpha}(h)=|\log h|$ and $02-\min\{\alpha-d/4;2\}=$ 0, and therefore $\epsilon_{h}\lesssim 1+|\log(h)|^{(d-3)/2}\lesssim 1$. Finally, for $\alpha-d/4>2$, $C_{\alpha}(h)=1$, $K_{\alpha}(h)=|\log h|^{(d-1)/2}$ and $02-\min\{\alpha-d/4;2\}=$ 0, and therefore $\epsilon_{h}\lesssim|\log h|^{-(d-1)/2}+1\lesssim 1$. In conclusion, we obtain for any $\alpha>d/4$, $\epsilon_{h}\lesssim 1$ which in turn yields
$\|\mathcal{Z}-\mathsf{Z}_{h}^{\ell}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}\lesssim K_{\alpha}(h)h^{\min\{\alpha-d/4;2\}}.\qed$
@@
3.4. Error analysis of the Chebyshev approximation
To bound the error between $\mathsf{Z}_{h,M}$ and $\mathsf{Z}_{h}$, we can directly apply Lang and Pereira (2023, Theorem 5.8). We obtain the following result.
Theorem 3.7.
@@ 3.4. Error analysis of the Chebyshev approximation
(39) $\|\mathsf{Z}_{h,M}-\mathsf{Z}_{h}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}_{h}))}\leq C_{V}^{-1}{h}^{-(d/2+1)}\exp(-C_{V}h^{-1}M)$
@@
Let $E_{\xi}\subset\amsmathbb{C}$ be the ellipse centered at $z=(\lambda_{\min}+\lambda_{\max})/2$, with foci $z_{1}=\lambda_{\min}$ and $z_{2}=\lambda_{\max}$, and semi-major axis $a_{\xi}=\lambda_{\max}/2$. In particular, note that $E_{\xi}\subset H_{\pi/2}$ and for any $z\in E_{\xi}$, $\operatorname{Re}(z)\geq\lambda_{\min}/2>0$. Hence, since $\gamma$ is a power spectral density, by definition $\gamma$ is holomorphic and bounded inside $E_{\xi}$. We can then adapt the same proof as in Lang and Pereira (2023, Theorem 5.8) to obtain the stated proposition. ∎
@@
For a fixed mesh size $h$, the approximation error converges to 00 as the order of the polynomial approximation $M$ goes to infinity. Choosing $M$ as a function of $h$ that grows fast enough then allows us to ensure the convergence of the approximation error as $n$ goes to infinity (Lang and Pereira, 2023, Section 5.2).
@@ Appendix A Deterministic proofs
A standard result in the spectral theory of elliptic operators on compact Riemannian manifolds (see e.g. (Shubin, 2001, Section 8)) ensures that there exists a set of eigenpairs $\{(\lambda_{i},f_{i})\}_{i\in\amsmathbb{N}}$ of $\mathcal{L}$ such that $0\leq\lambda_{1}\leq\lambda_{2}\leq\cdots$ with $\lambda_{i}\rightarrow+\infty$ as $i\rightarrow+\infty$, and $\{f_{i}\}_{i\in\amsmathbb{N}}$ is an orthonormal basis of $L^{2}(\mathcal{M})$ composed of possibly complex-valued functions.
@@
On the other hand, let $\lambda>0$ be one of the eigenvalues of $\mathcal{L}$, and $E_{\lambda}\subset L^{2}(\mathcal{M})$ the associated eigenspace. Following again the results from (Shubin, 2001, Section 8), we get that $E_{\lambda}\subset C^{\infty}(\mathcal{M})$, $\dim E_{\lambda}<\infty$ and that if $\lambda^{\prime}\neq\lambda$ is another eigenvalue of $\mathcal{L}$, then $E_{\lambda}$ and $E_{\lambda^{\prime}}$ are orthogonal. Besides, $E_{\lambda}$ is in fact generated by the set $\{f_{j}\}_{j\in J_{\lambda}}$, where $J_{\lambda}=\{i\in\amsmathbb{N}:\lambda_{i}=\lambda\}$ is finite (since $\dim E_{\lambda}<\infty$).
@@
A.2. Proof of Lemma 2.4
@@
Let us start by proving the following claims: for any $z\in\Gamma$, $f\in L^{2}(\mathcal{M})$,
(40) $\|(z-\mathcal{L}_{h})^{-1}P_{h}f-(z-\mathcal{L})^{-1}f\|_{L^{2}(\mathcal{M})}\leq Ch^{2},$
(41) $\|(z-\mathcal{L}_{h})^{-1}P_{h}f-(z-\mathcal{L})^{-1}f\|_{L^{2}(\mathcal{M})}\leq\frac{C}{|z|}.$
We omit the proof of Equation 41 and refer the reader to Fujita and Suzuki (1991, Theorem 5.1), where the proof may be adapted verbatim.
To prove Equation 40, we can largely proceed as in Fujita and Suzuki (1991, Theorem 7.1), substituting the Euclidean elliptic regularity estimates and interpolation bounds with their surface counterparts, see for instance Taylor (2011, Chapter 11), Dziuk and Elliott (2013, Lemma 4.3) and Bonito et al. (2024, Equation (4.8)). We must in essence only adapt Fujita and Suzuki (1991, Lemma 7.1) to also hold for $z\in\Gamma_{0}$, as their estimate only holds for $z\in\partial G_{\theta}$. In other words, we must show that for any $z=\delta_{0}e^{it}$, where $t\in[-\theta,\theta]$, there is a constant $\tilde{\delta}>0$ such that for all $v\in H^{1}(\mathcal{M})$,
(42) $|z|\|v\|_{L^{2}(\mathcal{M})}^{2}+\|v\|_{H^{1}(\mathcal{M})}^{2}\leq\tilde{\delta}^{-1}\left|z\|v\|_{L^{2}(\mathcal{M})}^{2}-\mathsf{A}_{\mathcal{M}}(v,v)\right|.$
To prove this inequality, note that by coercivity of $\mathsf{A}_{\mathcal{M}}$,
$\mathsf{A}_{\mathcal{M}}(v,v)-\delta_{0}\|v\|_{L^{2}(\mathcal{M})}^{2}\geq\delta\|v\|_{H^{1}(\mathcal{M})}^{2}-\delta_{0}\|v\|_{L^{2}(\mathcal{M})}^{2}.$
Let then $\tilde{\delta}=\frac{\delta_{0}}{1+\delta_{0}}$. Since the only requirement on $\delta_{0}$ was that $\delta_{0}<\delta/2$, we can then write
$\mathsf{A}_{\mathcal{M}}(v,v)-\delta_{0}\|v\|_{L^{2}(\mathcal{M})}^{2}$ $\geq 2\delta_{0}\|v\|_{H^{1}(\mathcal{M})}^{2}-\delta_{0}\|v\|_{L^{2}(\mathcal{M})}^{2}=2\tilde{\delta}(1+\delta_{0})\|v\|_{H^{1}(\mathcal{M})}^{2}-\tilde{\delta}(1+\delta_{0})\|v\|_{L^{2}(\mathcal{M})}^{2}$
$=2\tilde{\delta}(1+\delta_{0})\|v\|_{H^{1}(\mathcal{M})}^{2}+\left(\delta_{0}\tilde{\delta}-(2\delta_{0}+1)\tilde{\delta}\right)\|v\|_{L^{2}(\mathcal{M})}^{2}$
$\geq 2\tilde{\delta}(1+\delta_{0})\|v\|_{H^{1}(\mathcal{M})}^{2}-(2\delta_{0}+1)\tilde{\delta}\|v\|_{H^{1}(\mathcal{M})}^{2}+\delta_{0}\tilde{\delta}\|v\|_{L^{2}(\mathcal{M})}^{2}$
$=\tilde{\delta}\left(\|v\|_{H^{1}(\mathcal{M})}^{2}+\delta_{0}\|v\|_{L^{2}(\mathcal{M})}^{2}\right)=\tilde{\delta}\left(\|v\|_{H^{1}(\mathcal{M})}^{2}+|z|\|v\|_{L^{2}(\mathcal{M})}^{2}\right),$
since $|z|=\delta_{0}$ if $z\in\Gamma_{0}$. Therefore,
$\|v\|_{H^{1}(\mathcal{M})}^{2}+|z|\|v\|_{L^{2}(\mathcal{M})}^{2}\leq\tilde{\delta}^{-1}\left(\mathsf{A}_{\mathcal{M}}(v,v)-\delta_{0}\|v\|_{L^{2}(\mathcal{M})}^{2}\right).$
This last inequality implies in particular that $\mathsf{A}_{\mathcal{M}}(v,v)-\delta_{0}\|v\|_{L^{2}(\mathcal{M})}^{2}\geq 0$ and therefore that
$\|v\|_{H^{1}(\mathcal{M})}^{2}+|z|\|v\|_{L^{2}(\mathcal{M})}^{2}\leq\tilde{\delta}^{-1}\left|\mathsf{A}_{\mathcal{M}}(v,v)-\delta_{0}\|v\|_{L^{2}(\mathcal{M})}^{2}\right|.$
Using once again that $|z|=\delta_{0}$ and that $\mathsf{A}_{\mathcal{M}}(v,v)\geq 0$, we deduce that
$\|v\|_{H^{1}(\mathcal{M})}^{2}+|z|\|v\|_{L^{2}(\mathcal{M})}^{2}$ $\leq\tilde{\delta}^{-1}\left||\mathsf{A}_{\mathcal{M}}(v,v)|-|z|\,\|v\|_{L^{2}(\mathcal{M})}^{2}\right|\leq\tilde{\delta}^{-1}\left|\mathsf{A}_{\mathcal{M}}(v,v)-z\,\|v\|_{L^{2}(\mathcal{M})}^{2}\right|$
by the reverse triangle inequality. To conclude, we have proven that the relation (42) holds and equipped with this estimate, we can proceed as in the proof of Fujita and Suzuki (1991, Theorem 7.1) and obtain Equation 40.
Finally, by interpolating between the bounds obtained in Equations 40 and 41, we obtain that, for any $\beta\in[0,1]$,
$\|(z-\mathcal{L}_{h})^{-1}P_{h}f-(z-\mathcal{L})^{-1}f\|_{L^{2}(\mathcal{M})}\leq C\frac{h^{2\beta}}{|z|^{1-\beta}}.\qed$
Appendix B Error between functions of discrete operators
The aim of this section is to prove Equation 24. To do so, we rely on a geometric consistency estimate, and on extensions of the results of Bonito et al. (2024, Lemma A.1), which we start by stating and proving.

### v2

@@ 3.3. Error analysis of the SFEM discretization
$\|\mathcal{Z}-\mathsf{Z}_{h}^{\ell}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}\lesssim C_{\alpha}(h)h^{2\min\{\alpha-d/4;1\}}+|\log(h)|^{(d-1)/2}\leavevmode\nobreak\ h^{2}+h^{2}.$
Noe then that $2\min\{\alpha-d/4;1\}\leq 2$ and that $|\log(h)|^{(d-1)/2}\lesssim C_{\alpha}(h)$. Hence, we have the bound $|\log(h)|^{(d-1)/2}\leavevmode\nobreak\ h^{2}\lesssim C_{\alpha}(h)h^{2\min\{\alpha-d/4;1\}}$ which allows us to conclude that
$\|\mathcal{Z}-\mathsf{Z}_{h}^{\ell}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}\lesssim C_{\alpha}(h)h^{2\min\{\alpha-d/4;1\}},$
and the proof is complete. ∎
@@
3.4. Error analysis of the Galerkin–Chebyshev approximation
The error between the field $\mathcal{Z}$ and its Galerkin–Chebyshev approximation $\mathsf{Z}_{h,M}$ described in Section 3.2 is derived by combining Theorem 3.1 with an error bound between the SFEM–Galerkin approximation $\mathsf{Z}_{h}$ and the Galerkin–Chebyshev approximation $\mathsf{Z}_{h,M}$ obtained using [19, Theorem 5.8]. The latter bound is expressed in the next result.
Lemma 3.7.
@@ 3.4. Error analysis of the Galerkin–Chebyshev approximation
(44) $\|\mathsf{Z}_{h,M}-\mathsf{Z}_{h}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}_{h}))}\leq C_{V}^{-1}{h}^{-(d/2+1)}\exp(-C_{V}hM)$
@@
Let $E_{\xi}\subset\amsmathbb{C}$ be the ellipse centered at $z=(\lambda_{\min}+\lambda_{\max})/2$, with foci $z_{1}=\lambda_{\min}$ and $z_{2}=\lambda_{\max}$, and semi-major axis $a_{\xi}=\lambda_{\max}/2$. In particular, note that $E_{\xi}\subset H_{\pi/2}$ and for any $z\in E_{\xi}$, $\operatorname{Re}(z)\geq\lambda_{\min}/2>0$. Hence, since $\gamma$ is an amplitude spectral density, by definition $\gamma$ is holomorphic and bounded inside $E_{\xi}$. We can then adapt the same proof as in [19, Theorem 5.8] to obtain the stated proposition. ∎
@@
Hence, for a fixed mesh size $h$, the approximation error $\|\mathsf{Z}_{h,M}-\mathsf{Z}_{h}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}_{h}))}$ converges to 00 as the order of the polynomial approximation $M$ goes to infinity. Choosing $M$ as a function of $h$ that grows fast enough then allows us to ensure the convergence of the approximation error as $n$ goes to infinity [19, Section 5.2]. In particular, by taking $M=M_{h}=(d/2+3)C_{V}^{-1}h^{-1}|\log h|=(d/2+3)C_{V}^{-1}h^{-1}\log(h^{-1})$ the bound (44) becomes
(45) $\|\mathsf{Z}_{h,M}-\mathsf{Z}_{h}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}_{h}))}$ $\leq C_{V}^{-1}\exp((d/2+1)\log(h^{-1})-(d/2+3)\log(h^{-1}))$
$=C_{V}^{-1}h^{2}.$
Putting together Theorem 3.1 and Equation 45, we can now state the following result which provides a bound between the field $\mathcal{Z}$ and its Galerkin–Chebyshev approximation $\mathsf{Z}_{h,M_{h}}$.
Theorem 3.8.
Let $\gamma$ be an $\alpha$-amplitude spectral density with $\alpha>d/4$, let $h\in(0,h_{0})$ and take $M_{h}=(d/2+3)C_{V}^{-1}h^{-1}|\log h|$. Then, The strong error between the random field $\mathcal{Z}$ by its Galerkin–Chebyshev approximation $\mathsf{Z}_{h,M_{h}}^{\ell}$ satisfies the bound
(46) $\|\mathcal{Z}-\mathsf{Z}_{h,M_{h}}^{\ell}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}\lesssim C_{\alpha}(h)h^{2\min\{\alpha-d/4;1\}},$
where $C_{\alpha}(h)=|\log h|$ if $d/4<\alpha\leq 1$, $C_{\alpha}(h)=|\log h|^{3/2}$ if $21<\alpha<1+d/$ 2, and $C_{\alpha}(h)=|\log h|^{1/2}$ if $\alpha\geq 1+d/2$.
Remark 3.9.
In practice, it is common to “chop” a Chebyshev polynomial approximation, i.e. to truncate the polynomial approximation at an early order. Indeed, let us assume that we have computed the coefficients $\{c_{k}\}_{0\leq k\leq M_{h}}$ of the Chebyshev series up to the order $M=M_{h}$ and let $c_{\max}=\max_{0\leq k\leq M_{h}}|c_{k}|$. Let $\varepsilon>0$ be fixed but arbitrary and assume that there exists $m_{\varepsilon}\in\{0,\dots,M_{h}-1\}$ such that for any $k>m_{\varepsilon}$, $|c_{k}|/c_{\max}<\varepsilon$. The error between the field $\mathsf{Z}_{h,M_{h}}$ and its “chopped” counterpart $\mathsf{Z}_{h,m_{\varepsilon}}$ is given by
$\|\mathsf{Z}_{h,M_{h}}-\mathsf{Z}_{h,m_{\varepsilon}}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}_{h}))}^{2}$ $=\|\sum_{i=1}^{N_{h}}(P_{\gamma,M_{h}}(\Lambda_{i}^{h})-P_{\gamma,m_{\varepsilon}}(\Lambda_{i}^{h}))W_{i}^{h}E_{i}^{h}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}_{h}))}^{2}$
$=\sum_{i=1}^{N_{h}}(P_{\gamma,M_{h}}(\Lambda_{i}^{h})-P_{\gamma,m_{\varepsilon}}(\Lambda_{i}^{h}))^{2}$
where $P_{\gamma,M_{h}}(\lambda)=\sum_{k=0}^{M_{h}}c_{k}T_{k}(\lambda)$ is the $M_{h}$-th order Chebyshev approximation of $\gamma$ and $T_{k}$ denotes the $k$-th (shifted) Chebyshev polynomial. In particular, we have $|P_{\gamma,M_{h}}(\Lambda_{i}^{h})-P_{\gamma,m_{\varepsilon}}(\Lambda_{i}^{h})|=|\sum_{k=m_{\varepsilon}+1}^{M_{h}}c_{k}T_{k}(\lambda)|\leq\sum_{k=m_{\varepsilon}+1}^{M_{h}}|c_{k}|\leavevmode\nobreak\ |T_{k}(\lambda)|\leq\sum_{k=m_{\varepsilon}+1}^{M_{h}}|c_{k}|\leq(M_{h}-m_{\varepsilon})c_{\max}\varepsilon\leq M_{h}c_{\max}\varepsilon$. Hence, the error between $\mathsf{Z}_{h,M_{h}}$ and its “chopped” counterpart $\mathsf{Z}_{h,m_{\varepsilon}}$ satisfies
$\|\mathsf{Z}_{h,M_{h}}-\mathsf{Z}_{h,m_{\varepsilon}}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}_{h}))}$ $\leq N_{h}^{1/2}M_{h}c_{\max}\varepsilon\lesssim h^{-(1+d/2)}|\log h|\leavevmode\nobreak\ \varepsilon$
Hence, if we have $\varepsilon\lesssim|\log h|^{-1}h^{(3+d/2)}$ then we can conclude that $\|\mathsf{Z}_{h,M_{h}}-\mathsf{Z}_{h,m_{\varepsilon}}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}_{h}))}\lesssim h^{2}$, thus implying that once again we have $\|\mathcal{Z}-\mathsf{Z}_{h,m_{\varepsilon}}^{\ell}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}\lesssim C_{\alpha}(h)h^{2\min\{\alpha-d/4;1\}}$. This means that in practice, if we notice that the coefficients of the Chebyshev polynomial approximation $\{c_{k}\}_{0\leq k\leq M_{h}}$ decay fast enough that we can find some $m_{\varepsilon}\in\{0,\dots,M_{h}-1\}$ such that for any $k>m_{\varepsilon}$, $|c_{k}|/c_{\max}<\varepsilon\lesssim|\log h|^{-1}h^{(3+d/2)}$, we can replace the field $\mathsf{Z}_{h,M_{h}}$ by its chopped counterpart while still maintaining the same strong error bound with respect to the field $\mathcal{Z}$.
4. Numerical experiments
In this section, we present numerical experiments confirming the strong error bound in Equation 46. We consider two cases: when the manifold $\mathcal{M}$ is a circle (hence $d=1$), and when the manifold $\mathcal{M}$ is a sphere (hence $d=2$). Indeed, in both cases we are able to easily define nested meshes of various sizes $h$, and define appropriate noise projections at the different levels.
In each case, the field $\mathcal{Z}$ is approximated by a Galerkin–Chebyshev field $\mathsf{Z}_{h_{\text{fine}},M_{h_{\text{fine}}}}$ computed at a very fine mesh size $h_{\text{fine}}$. More precisely, 25 samples of $\{\mathsf{Z}_{h_{\text{fine}},M_{h_{\text{fine}}}}^{(i)}\}_{1\leq i\leq 25}$ of the field $\mathsf{Z}_{h_{\text{fine}},M_{h_{\text{fine}}}}$ are computed. In particular, let us denote by $\mathsf{W}^{(i)}_{h_{\text{fine}}}$ the projected white noise used to compute the sample $\mathcal{Z}^{(i)}=P_{\gamma,M_{h_{\text{fine}}}}(\mathsf{L}_{h_{\text{fine}}})\mathsf{W}^{(i)}_{h_{\text{fine}}}$. For a coarse mesh size $h>h_{\text{fine}}$, each sample $\mathcal{Z}^{(i)}$ is compared to the sample of $\mathsf{Z}_{h,M_{h}}^{(i)}$ computed from the white noise $\mathsf{W}^{(i)}_{h}$ obtained by projecting the noise $\mathsf{W}^{(i)}_{h_{\text{fine}}}$ (defined in the fine mesh) down to the coarse mesh. Note that, following Remark 3.9, the Galerkin–Chebyshev fields are systematically computed by chopping the Chebyshev polynomial approximations at a small level $\epsilon=10^{-12}$. Besides, when computing the representation of a Galerkin–Chebyshev field in the nodal basis through (36), the matrix $\sqrt{\bm{C}}$ is computed as a Cholesky factorization of the mass matrix $\bm{C}$ in the first experiment, and is approximate using a (diagonal) mass lumping approach in the second experiment. Finally, the strong error at a mesh size $h$ is then estimated from the samples using
(47) $\|\mathcal{Z}-\mathsf{Z}_{h,M_{h}}^{\ell}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}\approx\bigg{(}\frac{1}{25}\sum_{i=1}^{25}\|\mathsf{Z}_{h_{\text{fine}},M_{h_{\text{fine}}}}^{(i)}-\mathsf{Z}_{h,M_{h}}^{(i)}\|_{L^{2}(\mathcal{M}_{h})}^{2}\bigg{)}^{1/2}.$
In the first experiment, we consider a random field $\mathcal{Z}$ defined on the unit circle $\mathcal{M}$ with amplitude spectral density given by the formula $\gamma(\lambda)=(V_{0})^{\alpha}(\lambda+\cos(0.9\pi)\sqrt{\lambda})^{-\alpha}$ where $V_{0}=10^{4}$ and for $\alpha\in\{0.5,1.05,1.5\}$. The bilinear form $\mathcal{M}$ is taken as follows. The diffusion matrix $\mathcal{D}$ is defined by $\mathcal{D}(x)=I-\nu(x)\nu(x)^{T}$, $x\in\mathcal{M}$, where $\nu(x)=x/\|x\|$ is the normal vector to a point $x\in\mathcal{M}$ of the unit circle. The function $V$ is defined as $V(x)=3V_{0}$, $x\in\mathcal{M}$, if $\theta(x)\in]\pi/2,3\pi/2[$ and $V(x)=V_{0}$ otherwise, and $\theta(x)\in[0,2\pi[$ denotes the circular coordinate of the point $x\in\mathcal{M}$. The fine mesh size is taken to be $h_{\text{fine}}=2^{-19}\pi$ and the coarse mesh sizes are $h=\{2^{-13}\pi,2^{-14}\pi,2^{-15}\pi,2^{-16}\pi\}$. An example of sample (on the fine mesh) obtained with this choice of parameters is displayed in Figure 5(a). The results of this first experiment are presented in Figure 5(b), where the log of the strong error (47) is plotted against the log of the mesh size $h$, for the three amplitude spectral density obtained by taking $\alpha\in\{0.5,1.05,1.5\}$. As shown in the figure, we retrieve the rates (respectively $0.5$,$1.6$ and $2$) predicted by Theorem 3.8.
(a) Sample (in green) of the random field with $\alpha=1.75$ on the fine circle mesh (in black). The distance to the circle indicates the value of taken by the field.
(b) Strong error as a function the mesh size.
Figure 5. First numerical experiment (on the circle).
In the second experiment, we consider a random field $\mathcal{Z}$ defined on the unit sphere $\mathcal{M}$ with amplitude spectral density given by the formula $\gamma(\lambda)=C_{0}\lambda^{-\alpha}$ where $C_{0}=500$ and for $\alpha\in\{0.75,1.25,2.25\}$. The bilinear form $\mathcal{M}$ is taken as follows. The diffusion matrix $\mathcal{D}$ is defined $\mathcal{D}(x)=\nabla_{\mathcal{M}}f(x)(\nabla_{\mathcal{M}}f(x))^{T}+\rho(x)\nabla_{\mathcal{M}}^{\perp}f(x)(\nabla_{\mathcal{M}}^{\perp}f(x))^{T}$, where $f(x)=2\cos(\theta(x))\cos(\phi(x))\sin^{2}(\theta(x))$ and $\rho(x)=0.1+0.6/(1+e^{-4\cos(\theta(x))})$. Here $x\in\mathcal{M}$ and $(\theta(x),\phi(x))$ denotes spherical coordinates. The function $V$ is defined as $V(x)=500(1+5\cos^{2}(\pi\theta(x)))$. The fine mesh size is taken to be $h_{\text{fine}}=2.16\times 10^{-3}$ and the coarse mesh sizes are $h=\{3.45\times 10^{-2},1.73\times 10^{-2},8.63\times 10^{-3},4.32\times 10^{-3}\}$. An example of sample (on the fine mesh) obtained with this choice of parameters is displayed in Figure 6(a). The results of this second experiment are presented in Figure 6(b), where once again the log of the strong error (47) is plotted against the log of the mesh size $h$, for the three amplitude spectral density obtained by taking $\alpha\in\{0.75,1.25,2.25\}$. As seen in Figure 6(b), we retrieve the rates (respectively $0.5$,$1.5$ and $2$) predicted by Theorem 3.8.
(a) Sample of the random field with $\alpha=1.75$ on the fine mesh. The colors indicate the value of taken by the field.
(b) Strong error as a function the mesh size.
Figure 6. Second numerical experiment (on the sphere).
@@ Appendix A Deterministic proofs
A standard result in the spectral theory of elliptic operators on compact Riemannian manifolds (see e.g. [23, Section 8]) ensures that there exists a set of eigenpairs $\{(\lambda_{i},f_{i})\}_{i\in\amsmathbb{N}}$ of $\mathcal{L}$ such that $0\leq\lambda_{1}\leq\lambda_{2}\leq\cdots$ with $\lambda_{i}\rightarrow+\infty$ as $i\rightarrow+\infty$, and $\{f_{i}\}_{i\in\amsmathbb{N}}$ is an orthonormal basis of $L^{2}(\mathcal{M})$ composed of possibly complex-valued functions.
@@
On the other hand, let $\lambda>0$ be one of the eigenvalues of $\mathcal{L}$, and $E_{\lambda}\subset L^{2}(\mathcal{M})$ the associated eigenspace. Following again the results from [23, Section 8], we get that $E_{\lambda}\subset C^{\infty}(\mathcal{M})$, $\dim E_{\lambda}<\infty$ and that if $\lambda^{\prime}\neq\lambda$ is another eigenvalue of $\mathcal{L}$, then $E_{\lambda}$ and $E_{\lambda^{\prime}}$ are orthogonal. Besides, $E_{\lambda}$ is in fact generated by the set $\{f_{j}\}_{j\in J_{\lambda}}$, where $J_{\lambda}=\{i\in\amsmathbb{N}:\lambda_{i}=\lambda\}$ is finite (since $\dim E_{\lambda}<\infty$).
@@
A.2. A useful inequality
We end this section by introducing a lemma which will be used to obtain practical error bounds.
Lemma A.1.
For any $a,b\in\amsmathbb{R}$ and any $\varepsilon>0$, $\min\{a-\varepsilon;b\}+\varepsilon\geq\min\{a;b\}$.
@@ A.2. A useful inequality
Indeed, if $a<b$, we have $\min\{a-\varepsilon;b\}+\varepsilon=a-\varepsilon+\varepsilon=a=\min\{a;b\}$. If $b\leq a<b+\varepsilon$, $\min\{a-\varepsilon;b\}+\varepsilon=a-\varepsilon+\varepsilon=a\geq b=\min\{a;b\}$. And finally if $a\geq b+\varepsilon$, $\min\{a-\varepsilon;b\}+\varepsilon=b+\varepsilon>b=\min\{a;b\}$. ∎
Appendix B Error estimates